OpenStudy (anonymous):

Multiply. (3sqrt3 -6sqrt6)(2sqrt3 +4sqrt6)=

OpenStudy (anonymous):

\[(3 \sqrt3 -6\sqrt6)(2\sqrt3 +4\sqrt6)\]

OpenStudy (anonymous):

= 3*2*3-12*3root2 +12*3root2 -24*6= 18-144= -126